OpenStudy (anonymous):

How do I solve this: 2(cos^2 x - 1) = -1

OpenStudy (anonymous):

\[x=(\frac{ \pi n }{ 2 }-\frac{ \pi }{ 4 } ) , n \in z\]

OpenStudy (anonymous):

Can you please explain how you solved this? I have a test tommorow and need to be able to repeat these procedures.

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

Solve for x: 2 (cos^2(x)-1) = -1 Divide both sides by a constant to simplify the equation. Divide both sides by 2: cos^2(x)-1 = -1/2 Isolate terms with x to the left hand side. Add 1 to both sides: cos^2(x) = 1/2 Eliminate the exponent on the left hand side. Take the square root of both sides: cos(x) = 1/sqrt(2) or cos(x) = -1/sqrt(2) Look at the first equation: Eliminate the cosine from the left hand side. Take the inverse cosine of both sides: x = pi/4+2 pi n_1 for n_1 element Z or x = (7 pi)/4+2 pi n_2 for n_2 element Z or cos(x) = -1/sqrt(2) Look at the third equation: Eliminate the cosine from the left hand side. Take the inverse cosine of both sides: Answer: | | x = pi/4+2 pi n_1 for n_1 element Z or x = (7 pi)/4+2 pi n_2 for n_2 element Z or x = (3 pi)/4+2 pi n_3 for n_3 element Z or x = (5 pi)/4+2 pi n_4 for n_4 element Z